OpenStudy (anonymous):

please helpppp Is the expression x3•x3•x3 equivalent to x3•3•3? Why or why not? Explain your reasoning

OpenStudy (anonymous):

\[x^3*x^3*x^3=x^(3+3+3)=x^9\]

OpenStudy (anonymous):

the seond one should be x to the power of 3 times 3 times 3

OpenStudy (anonymous):

and \[x^3*3*3=(9)(x^3)=9x^3\]

OpenStudy (anonymous):

For 1. The exponential law states that when you have an integer x to some number n in the form \[x^n\] and another same integer x to some number y in the form \[x^y\], then you can multiply the coffefficients and add the exponents to make \[x^(n+y)\]

OpenStudy (anonymous):

for the second one due to your edited question, just multiply 9 by 3 to get \[27x^3\]

OpenStudy (anonymous):